# Autoencoding Random Forests

### v1

@@ 6 Discussion
One notable difference between our method and autoencoding neural networks is that RFAE is not trained end-to-end. That is, while a deep autoencoder simultaneously learns to encode and decode, RFAE is effectively a post-processing procedure for a pre-trained RF, with independent modules for encoding and decoding. We highlight that end-to-end training represents a fundamentally different objective. Whereas traditional autoencoders are necessarily unsupervised, our method works in tandem with either supervised or unsupervised RFs. As such, the goal is not necessarily to learn an efficient representation for its own sake, but rather to reveal the inner workings of a target model. One upshot of this decoupling is that our split relabeling and $k$-NN decoders can work in tandem with any valid encoding scheme. For instance, we could relabel an RF’s splits to approximate the behavior of sample points in principal component space, or indeed any $\mathcal{Z}$ for which we have a map $g:\mathcal{X}\mapsto\mathcal{Z}$.
We highlight two notable limitations of our approach. First, the computational demands of our decoding strategies are nontrivial. (For a detailed analysis, see Appx. D.) Second, when using the $k$-NN approach, results will vary with the choice of $k$. However, we observe that autoencoding is a difficult task in general, and top deep learning models pose far greater computational burdens than RFAE. Moreover, having just a single hyperparameter to worry about is a rare luxury in this field, where leading algorithms often require bespoke architectures and finely tuned regularization penalties. Compared to the leading alternatives, RFAEs are relatively lightweight and simple.
@@ Appendix A Proofs
Under (A1)-(A5), decision trees satisfy the criteria of Stone’s theorem [84] and are therefore universally consistent (see Devroye et al. [28, Thm. 6.1] and Györfi et al. [45, Thm. 4.2]). The consistency of the ensemble follows from the consistency of the basis functions [12]. There is some debate in the literature as to whether these assumptions are necessary for universal consistency—(A1) and (A2) in particular may be overly strong—but they are provably sufficient. See Biau [10, Rmk. 8], Wager and Athey [93, Appx. B], and Tang et al. [85] for a discussion.
@@
    By regularity conditions (A3) and (A4), every coordinate has a nonzero probability $\rho>0$ of being selected for splitting at any node, and every split allocates at least a fraction $\gamma\in(0,0.5]$ of the points to each child node. Meinshausen [65, Lemma 2] has shown that, under these conditions, the diameter of each leaf goes to zero in probability. This amount to an asymptotic injectivity guarantee. Since $\mathcal{X}$ is compact, for any pair of distinct points $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$, there exists an index $\ell$ such that some $f_{\ell}\in\mathcal{H}_{\ell}$ assigns different values to $\mathbf{x}$ and $\mathbf{x}^{\prime}$. In other words, some tree in $\mathcal{A}$ is guaranteed to route $\mathbf{x}$ and $\mathbf{x}^{\prime}$ to different leaves.
@@
Item (c) follows from (b) under our definition of universality. This was first shown by Gretton et al. [40, Thm. 3] in the non-asymptotic regime (although Sriperumbudur et al. [82] prove that the properties can come apart under slightly different notions of universality). We adapt the result simply by substituting a sufficiently close approximation in $\mathcal{H}$, where proximity is defined w.r.t. the supremum norm. Since the existence of such a close approximation is guaranteed by (b), the sequence {$\{\mathcal{H}_{\ell}\}$ is asymptotically characteristic.
@@
This result follows trivially from the universal consistency of the RF algorithm itself. Any partition of $\mathcal{X}$ that satisfies regularity conditions (A1)-(A5) converges on the true joint distribution $P_{X}$ as $n,d_{\Phi}\rightarrow\infty$ (see [63, Thm. 1] and [23, Thm. 2]). Resulting leaves have shrinking volume, effectively converging on individual points $\mathbf{x}$ with coverage proportional to the density $p(\mathbf{x})$. Because there are no gaps in these leaves (i.e., no subregions of zero density), a weighted mixture of uniform draws—with weights given by the leaf coverage—is asymptotically equivalent to sampling from $P_{X}$ itself. A leaf assignment oracle would therefore be guaranteed to map each latent vector $\mathbf{z}\in\mathcal{Z}$ to the corresponding input point $\mathbf{x}\in\mathcal{X}$, since leaf assignments effectively determine feature values in the large sample limit.
@@
Such counterexamples become increasingly rare as the forest grows more complex. As previously noted, under (A3) and (A4), leaf diameter vanishes in probability [65, Lemma 2]. This is why random forests are asymptotically injective, a property we exploited in the proof for part (b) of Thm. 3.4. A corollary of Meinshausen’s lemma is that for all distinct points $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$:
@@
[…]
@@
[…]
@@ B.1 Reconstruction benchmark
[…]
@@
  • Autoencoder: We use an MLP-based autoencoder for this benchmark. This is in contrast to a CNN-based autoencoder, which is more suited to image data, but is not relevant in tabular data tasks, where there is no inherent structure that could be extracted by convolution kernels. We structure our network to have five hidden layers, where the size of these layers are adaptive to $d_{X}$ and $d_{\mathcal{Z}}$. In particular, we want to structure the network such that the size of each hidden layer reduces uniformly from $d_{X}$ to $d_{\mathcal{Z}}$ at encoding and increases uniformly from $d_{\mathcal{Z}}$ to $d_{X}$ at decoding. Our structure then is:
    Input($d_{X}$) $\rightarrow$ Dense($d_{X}-(d_{X}-d_{\mathcal{Z}})\times 1/3$) $\rightarrow$ Dense($d_{X}-(d_{\mathcal{Z}}-d_{\mathcal{Z}})\times 2/3$) $\rightarrow$ Latent($d_{\mathcal{Z}}d_{\mathcal{Z}}$) $\rightarrow$ Dense($d_{X}-(d_{X}-d_{\mathcal{Z}})\times 2/3$) $\rightarrow$ Dense($d_{X}-(d_{X}-d_{\mathcal{Z}})\times 1/3$) $\rightarrow$ Output($d_{X}$).
    If $d_{X}=8,d_{\mathcal{Z}}=2$ then with this rule the network will be $8\rightarrow 6\rightarrow 4\rightarrow 2\rightarrow 4\rightarrow 6\rightarrow 8$.
@@
  • Variational Autoencoder: Similar to the autoencoder, we use an MLP-based variational autoencoder. For comparison, we mimic the autoencoder’s architecture, activation function, and defaults, only changing epochs = 100 and adding batch_size = 32. We also impose a $\beta$ coefficient inspired from [48]. However, we select a $\beta=0.1$, to avoid pos
@@
We present the results in Table 2 corresponding to the plots in Fig. 4.
Table 2: Mean distortion (with standard deviation) for each method and dataset, across all $d_{\mathcal{Z}}$ values used and all runs. Best result per row is bolded.
Dataset RFAE TVAE TTVAE AE VAE
abalone 0.167 (0.002) 0.309 (0.005) 0.260 (0.003) 0.230 (0.025) 0.211 (0.006)
adult 0.326 (0.007) 0.158 (0.005) 0.195 (0.007) 0.401 (0.003) 0.391 (0.004)
banknote 0.100 (0.012) 0.312 (0.013) 0.276 (0.023) 0.724 (0.023) 0.771 (0.013)
bc 0.333 (0.003) 0.564 (0.003) 0.359 (0.005) 0.287 (0.008) 0.578 (0.003)
car 0.320 (0.011) 0.195 (0.014) 0.107 (0.015) 0.349 (0.012) 0.313 (0.011)
churn 0.352 (0.012) 0.603 (0.011) 0.422 (0.014) 0.861 (0.005) 0.731 (0.006)
credit 0.315 (0.004) 0.450 (0.005) 0.375 (0.011) 0.450 (0.005) 0.456 (0.004)
diabetes 0.479 (0.016) 0.726 (0.007) 0.643 (0.014) 0.799 (0.011) 0.895 (0.004)
dry_bean 0.137 (0.002) 0.273 (0.002) 0.303 (0.008) 0.083 (0.014) 0.206 (0.001)
hd 0.432 (0.008) 0.582 (0.003) 0.605 (0.006) 0.892 (0.003) 0.916 (0.002)
king 0.308 (0.008) 0.352 (0.006) 0.348 (0.008) 0.377 (0.011) 0.518 (0.004)
marketing 0.292 (0.009) 0.304 (0.005) 0.259 (0.011) 0.357 (0.007) 0.372 (0.004)
mushroom 0.083 (0.001) 0.093 (0.003) 0.011 (0.003) 0.055 (0.004) 0.035 (0.004)
obesity 0.227 (0.008) 0.354 (0.004) 0.299 (0.008) 0.306 (0.009) 0.358 (0.003)
plpn 0.176 (0.006) 0.282 (0.006) 0.224 (0.011) 0.384 (0.013) 0.410 (0.009)
spambase 0.558 (0.005) 0.825 (0.002) 0.807 (0.003) 0.446 (0.010) 0.784 (0.001)
student 0.371 (0.002) 0.424 (0.001) 0.426 (0.004) 0.536 (0.003) 0.551 (0.002)
telco 0.177 (0.003) 0.155 (0.003) 0.091 (0.007) 0.128 (0.005) 0.130 (0.005)
wq 0.240 (0.005) 0.691 (0.008) 0.759 (0.006) 0.467 (0.019) 0.708 (0.004)
Average Rank 1.80 3.38 2.45 3.27 4.10
@@ B.2 Additional results
We present the results of additional experiments comparing different decoding methods, multip as well as supervised vs. unsupervised RF embeddings.
@@
We present the results in figure 6, where the kNN method is denoted as RFAE.
@@
Figure 7 complements Figure 3 in Section 5 by showing the effects of different parameters on the reconstruction performance for MNIST test digits: the number of trees $B$ (Figure 7(a)), the number of samples $n_{\text{train\_enc}}$ used for encoder training (Figure 7(b)), the diffusion time step $t$ (Figure 7(c)), and the number $k$ of nearest neighbors used for decoding (Figure 7(d)). In each subplot, the respective parameter varies in a pre-defined range, while all other ones are kept fixed at $B=1000$, $n_{\text{train\_enc}}=30\,000$, $t=1$, and $k=50$; the latent dimension is set to $d_{\mathcal{Z}}=32$ throughout.
@@
Figure 8 compares RFAE reconstructions using supervised and (unsupervised) completely random forests with those produced by a convolutional autoencoder with three convolutional layers. All models were trained on full training data and with $d_{\mathcal{Z}}=32$.
@@
RF embeddings provide different perspectives on the data depending on whether we use supervised or unsupervised learning objectives. We train a standard RF classifier to distinguish acute lymphoblastic (ALL) from acute myeloid (AML) leukemia samples in a well known transcriptomic dataset [38]. After filtering, the feature matrix includes expression values for $d_{X}=3572$ genes recorded in $n=72$ patients. These kind of short, wide datasets are common in bioinformatics, but can be challenging for many learning algorithms. RFs excel in these settings, attaining out-of-bag accuracy of 98% on this task. KPC1 clearly separates the two classes in the left panel, while KPC2 appears to isolate a potential outlier within the AML cohort. Using an unsupervised adversarial RF [94], we find far greater overlap among the two classes (as expected), although ALL and AML samples are hardly uniform throughout the latent space. This example demonstrates the flexibility of our method. Whereas standard spectral embedding algorithms with fixed kernels are limited to unsupervised representations, RFAE can take any RF input.

### v2

@@ 6 Discussion
One notable difference between our method and autoencoding neural networks is that RFAE is not trained end-to-end. That is, while a deep autoencoder simultaneously learns to encode and decode, RFAE is effectively a post-processing procedure for a pre-trained RF, with independent modules for encoding and decoding. We highlight that end-to-end training represents a fundamentally different objective. Whereas traditional autoencoders are necessarily unsupervised, our method can be readily applied to regression or classification forests, linking RFAE to supervised dimensionality reduction techniques [38, 6]. More generally, our goal is not just to learn an efficient representation for its own sake, but rather to reveal the inner workings of a target model. One upshot of this decoupling is that our split relabeling and $k$-NN decoders can work in tandem with any valid encoding scheme. For instance, we could relabel an RF’s splits to approximate the behavior of sample points in principal component space, or indeed any $\mathcal{Z}$ for which we have a map $g:\mathcal{X}\mapsto\mathcal{Z}$.
We highlight two notable limitations of our approach. First, the computational demands of our decoding strategies are nontrivial. (For a detailed analysis, see Appx. D.) Second, when using the $k$-NN approach, results will vary with the choice of $k$. (Experimental results on hyperparameter sensitivity are presented in Appx. B.2.) However, we observe that autoencoding is a difficult task in general. Top deep learning models generally pose far greater computational burdens than RFAE, and require many more hyperparameters to govern model architecture and regularization penalties. Compared to the leading alternatives, RFAEs are relatively lightweight and user-friendly.
@@ Appendix A Proofs
Under (A1)-(A5), decision trees satisfy the criteria of Stone’s theorem [89] and are therefore universally consistent (see Devroye et al. [32, Thm. 6.1] and Györfi et al. [50, Thm. 4.2]). The consistency of the ensemble follows from the consistency of the basis functions [15]. There is some debate in the literature as to whether these assumptions are necessary for universal consistency—(A1) and (A2) in particular may be overly strong—but they are provably sufficient. See Biau [13, Rmk. 8], Wager and Athey [98, Appx. B], and Tang et al. [90] for a discussion.
@@
It may not be immediately obvious that Eq. 1 is equivalent to the scaled inner product $B^{-1}\big\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\big\rangle$. For completeness, we derive the identity. Expanding the inner product, we have:
$\big\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\big\rangle$ $=\sum_{b=1}^{B}\penalty 10000\ \big\langle\phi^{(b)}(\mathbf{x}),\phi^{(b)}(\mathbf{x}^{\prime})\big\rangle$
$=\sum_{b=1}^{B}\penalty 10000\ \Big\langle\pi^{(b)}(\mathbf{x})\odot\sqrt{\mathbf{s}^{(b)}},\pi^{(b)}(\mathbf{x}^{\prime})\odot\sqrt{\mathbf{s}^{(b)}}\Big\rangle$
$=\sum_{b=1}^{B}\sum_{\ell=1}^{L_{b}}\penalty 10000\ \pi_{\ell}^{(b)}(\mathbf{x})\penalty 10000\ \pi_{\ell}^{(b)}(\mathbf{x}^{\prime})\penalty 10000\ s_{\ell}^{(b)}.$
If both $\mathbf{x}$ and $\mathbf{x}^{\prime}$ fall in leaf $\ell$ of tree $b$, then $\pi_{\ell}^{(b)}(\mathbf{x})\penalty 10000\ \pi_{\ell}^{(b)}(\mathbf{x}^{\prime})=1$; otherwise, the product evaluates to 0. Let $s_{\text{leaf}}^{(b)}$ denote the inverse sample size of the leaf containing $\mathbf{x}$ in tree $b$. Then $\big\langle\phi^{(b)}(\mathbf{x}),\phi^{(b)}(\mathbf{x}^{\prime})\big\rangle=s_{\text{leaf}}^{(b)}$ if $\mathbf{x},\mathbf{x}^{\prime}$ fall in the same leaf of tree $b$, or 0 otherwise. Note that we have:
$s_{\text{leaf}}^{(b)}=\frac{1}{\sum_{i=1}^{n}k^{(b)}(\mathbf{x},\mathbf{x}_{i})},$
since the number of training points in the same leaf as $\mathbf{x}$ is given by $\sum_{i=1}^{n}k^{(b)}(\mathbf{x},\mathbf{x}_{i})$. Substituting and dividing by $B$ gives:
$\frac{1}{B}\big\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\big\rangle=\frac{1}{B}\sum_{b=1}^{B}\frac{k^{(b)}(\mathbf{x},\mathbf{x}^{\prime})}{\sum_{i=1}^{n}k^{(b)}(\mathbf{x},\mathbf{x}_{i})},$
which completes the derivation.
@@
    By regularity conditions (A3) and (A4), every coordinate has a nonzero probability $\rho>0$ of being selected for splitting at any node, and every split allocates at least a fraction $\gamma\in(0,0.5]$ of the points to each child node. Meinshausen [70, Lemma 2] has shown that, under these conditions, the diameter of each leaf goes to zero in probability. This amount to an asymptotic injectivity guarantee. Since $\mathcal{X}$ is compact, for any pair of distinct points $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$, there exists an index $\ell$ such that some $f_{\ell}\in\mathcal{H}_{\ell}$ assigns different values to $\mathbf{x}$ and $\mathbf{x}^{\prime}$. In other words, some tree in $\mathcal{A}$ is guaranteed to route $\mathbf{x}$ and $\mathbf{x}^{\prime}$ to different leaves.
@@
Item (c) follows from (b) under our definition of universality. This was first shown by Gretton et al. [45, Thm. 3] in the non-asymptotic regime (although Sriperumbudur et al. [87] prove that the properties can come apart under slightly different notions of universality). We adapt the result simply by substituting a sufficiently close approximation in $\mathcal{H}$, where proximity is defined w.r.t. the supremum norm. Since the existence of such a close approximation is guaranteed by (b), the sequence $\{\mathcal{H}_{\ell}\}$ is asymptotically characteristic.
@@
This result follows trivially from the universal consistency of the RF algorithm itself. Any partition of $\mathcal{X}$ that satisfies regularity conditions (A1)-(A5) converges on the true joint distribution $P_{X}$ as $n,d_{\Phi}\rightarrow\infty$ (see [68, Thm. 1] and [27, Thm. 2]). Resulting leaves have shrinking volume, effectively converging on individual points $\mathbf{x}$ with coverage proportional to the density $p(\mathbf{x})$. Because there are no gaps in these leaves (i.e., no subregions of zero density), a weighted mixture of uniform draws—with weights given by the leaf coverage—is asymptotically equivalent to sampling from $P_{X}$ itself. A leaf assignment oracle would therefore be guaranteed to map each latent vector $\mathbf{z}\in\mathcal{Z}$ to the corresponding input point $\mathbf{x}\in\mathcal{X}$, since leaf assignments effectively determine feature values in the large sample limit.
@@
Such counterexamples become increasingly rare as the forest grows more complex. As previously noted, under (A3) and (A4), leaf diameter vanishes in probability [70, Lemma 2]. This is why random forests are asymptotically injective, a property we exploited in the proof for part (b) of Thm. 3.4. A corollary of Meinshausen’s lemma is that for all distinct points $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}$:
@@
    $\tilde{\mathbf{x}}\overset{p}{\rightarrow}\mathbf{x}$: sampling uniformly from the intersection of each sample’s assigned leaf regions asymptotically recovers the original inputs; and
@@
    $\hat{\mathbf{x}}\overset{p}{\rightarrow}\tilde{\mathbf{x}}$: $k$-NN regression in the spectral embedding space is universally consistent.
@@ B.1 Reconstruction benchmark
  • For each dataset, we define 10 compression ratios, or latent rates, going uniformly from 0.1 (10% of the number of features) to 1 (100% the number of features). For each value, we find $d_{\mathcal{Z}}=d_{\mathcal{X}}\times\text{latent\_rate}$
@@
  • Autoencoder: We use an MLP-based autoencoder for this benchmark. This is in contrast to a CNN-based autoencoder, which is more suited to image data, but is not relevant in tabular data tasks, where there is no inherent structure that could be extracted by convolution kernels. We structure our network to have five hidden layers, where the size of these layers are adaptive to $d_{\mathcal{X}}$ and $d_{\mathcal{Z}}$. In particular, we want to structure the network such that the size of each hidden layer reduces uniformly from $d_{\mathcal{X}}$ to $d_{\mathcal{Z}}$ at encoding and increases uniformly from $d_{\mathcal{Z}}$ to $d_{\mathcal{X}}$ at decoding. Our structure then is:
[…]
    If $d_{\mathcal{X}}=8,d_{\mathcal{Z}}=2$ then with this rule the network will be $8\rightarrow 6\rightarrow 4\rightarrow 2\rightarrow 4\rightarrow 6\rightarrow 8$.
@@
  • Variational Autoencoder: Similar to the autoencoder, we use an MLP-based variational autoencoder. For comparison, we mimic the autoencoder’s architecture, activation function, and defaults, only changing epochs = 100 and adding batch_size = 32. We also impose a $\beta$ coefficient inspired from [53], at a value of 0.1.
@@
We present the results in Fig. 5 corresponding to the plots in Table 1.
@@ B.2 Additional results
We present the results of additional experiments comparing different decoding methods, hyperparameter analysis on MNIST, supervised vs. unsupervised RF embeddings as well as runtime experiments.
@@
We present the results in Fig. 6, where the kNN method is denoted as RFAE.
@@
Fig. 7 complements Fig. 3 in Section 5 by showing the effects of different parameters on the reconstruction performance for MNIST test digits: the number of trees $B$ (Fig. 7(a)), the number of samples $n_{\text{train\_enc}}$ used for encoder training (Fig. 7(b)), the diffusion time step $t$ (Fig. 7(c)), and the number $k$ of nearest neighbors used for decoding (Fig. 7(d)). In each subplot, the respective parameter varies in a pre-defined range, while all other ones are kept fixed at $B=1000$, $n_{\text{train\_enc}}=30\,000$, $t=1$, and $k=50$; the latent dimension is set to $d_{\mathcal{Z}}=32$ throughout.
@@
Fig. 8 compares RFAE reconstructions using supervised and (unsupervised) completely random forests with those produced by a convolutional autoencoder with three convolutional layers. All models were trained on full training data and with $d_{\mathcal{Z}}=32$.
@@
RF embeddings provide different perspectives on the data depending on whether we use supervised or unsupervised learning objectives. We train a standard RF classifier to distinguish acute lymphoblastic (ALL) from acute myeloid (AML) leukemia samples in a well known transcriptomic dataset [43]. After filtering, the feature matrix includes expression values for $d_{\mathcal{X}}=3572$ genes recorded in $n=72$ patients. These kind of short, wide datasets are common in bioinformatics, but can be challenging for many learning algorithms. RFs excel in these settings, attaining out-of-bag accuracy of 98% on this task. KPC1 clearly separates the two classes in the left panel, while KPC2 appears to isolate a potential outlier within the AML cohort. Using an unsupervised adversarial RF [99], we find far greater overlap among the two classes (as expected), although ALL and AML samples are hardly uniform throughout the latent space. This example demonstrates the flexibility of our method. Whereas standard spectral embedding algorithms with fixed kernels are limited to unsupervised representations, RFAE can take any RF input.
Runtime Experiments
Our runtime experiments consists of two tests; one where we compare the runtime of methods within the reconstruction benchmark (see B.1 for method and dataset details), and one where we exclusively examine the runtime of RFAE more closely. For the first experiment, we conduct a controlled runtime comparison of all five methods (AE, VAE, TVAE, TTVAE and RFAE) on three small-to-medium datasets (plpn, student, credit). These experiments are conducted on a laptop with Intel Core i5-10300H CPU, NVIDIA GeForce GTX 1650 (4GB), and 24GB DDR4 RAM. Each method was run 100 times across varying compression ratios, according to the same specifications as in the benchmark. We report the runtimes in Table 3, in seconds.
Table 3: Runtime comparison across five models and three datasets.
Dataset AE VAE RFAE TVAE TTVAE
plpn 161.6 185.9 523.8 744.2 2252.8
student 190.9 227.8 1058.6 3611.2 5148.7
credit 223.2 277.3 1379.5 2212.5 4046.0
Here, the results indicate RFAE’s superiority in runtime over the other two state-of-the-art methods in TVAE and TTVAE, while being slower than the naive AE and VAEs - which did poorly in the actual benchmark.
For the second experiment, we evaluate RFAE alone across 20 datasets, using a fixed compression ratio of 0.2 (10 runs per dataset). These were run on a HPC unit with an AMD Threadripper 3960X (24 cores, 48 threads) CPU and 256GB RAM, with no GPU unit used for the experiment. We report the average runtime for training and inference (as well as the total time), and their corresponding standard errors, ordered by average total runtime, in Table 4, in seconds. The results suggest RFAE scales well with dataset size.
Table 4: RFAE runtime performance across datasets (sorted by total mean).
Dataset Training $\mu$ Training $\sigma$ Inference $\mu$ Inference $\sigma$ Total $\mu$ Total $\sigma$ Samples # Features
plpn 1.89 0.00 1.35 0.01 3.24 0.03 333 8
credit 2.82 0.32 3.18 0.01 6.01 0.41 1,000 21
car 2.90 0.00 2.87 0.00 5.77 0.00 1,728 7
student 3.03 0.41 3.30 0.03 6.33 0.63 649 33
diabetes 3.16 0.01 3.97 0.01 7.13 0.01 768 9
hd 4.17 0.00 5.99 0.00 10.16 0.00 298 14
forestfires 4.63 0.30 6.12 0.02 10.76 0.36 517 13
bc 6.88 1.11 2.65 0.01 9.53 1.17 570 32
abalone 6.82 0.02 5.58 0.03 12.39 0.03 4,177 9
wq 7.90 0.73 6.38 0.01 14.28 0.81 4,898 12
obesity 8.74 0.02 12.11 0.08 20.85 0.05 2,111 16
churn 12.28 0.18 11.75 0.08 24.03 0.24 10,000 11
mushroom 13.80 0.23 11.53 0.10 25.33 0.48 8,124 22
telco 13.96 0.01 16.45 0.05 30.41 0.09 7,032 20
spambase 29.52 0.90 27.78 0.51 57.30 1.12 4,601 59
dry_bean 43.19 5.09 19.63 0.10 62.82 5.25 13,611 17
king 54.31 42.78 32.04 1.87 86.36 44.36 21,613 19
adult 150.27 304.30 89.47 24.85 239.75 469.52 45,222 14
marketing 148.96 4831.41 91.50 374.82 240.46 7889.97 45,211 17